A model-train transformer plugs into 110 V ac and draws 0.29 A while supplying 7.6 A to the train. What voltage is present across the tracks? Is the transformer step-up or step-down?
Answer #1

here,

the volatge across primary , Vp = 110 V

the current across primary , Ip = 0.29 A

the current across secondary , Is = 7.6 A

let the voltage across secondary be Vs

for an ideal transformer

Vp * Ip = Vs * Is

110 * 0.29 = 7.6 * Vs

Vs = 4.2 V

the voltage present across the tracks is 4.2 V

as Vs < Vp

the transformer is Step-down